# Beyond Any-Shot Adaptation: Predicting Optimization Outcome for Robustness Gains without Extra Pay

### v2

@@ 3.3 MPTS still achieves adaptation robustness in zero-shot continuous control
Figure 6: DR Results on Ergo-Reacher and Lunar-Lander (7 Runs). a. In Ergo-Reacher, the $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average cumulative returns on validation MDPs are reported together with the risk learner’s PCC curve during DR training. b. In Lunar-Lander, the cumulative returns on validation MDPs are illustrated together with the risk learner’s PCC curve during DR training. c. We test the DR trained policies on the fixed MDP set and report the cumulative returns of the $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$ and $\text{CVaR}_{0.5}$ ones. d. The testing returns averaged over the fixed collection of MDPs are illustrated. e. The required runtime is computed for all methods on Lunar-Lander. f. In Lunar-Lander, shown are frequencies of sampled identifiers using MPTS during DR training. g. In Lunar-Lander, we test the trained policies in both in-distribution (ID) domains and out-of-distribution (OOD) domains to report each task’s average returns.
MPTS learns the least tasks but achieves the most in DR training.
The Ergo-Reacher curves contradict general meta-RL findings; the use of robust optimization, such as MPTS and DRM, improves both the average and CVaR performance in Fig. 6a. It seems that MPTS receives an extra task exploration bonus from the relatively larger $\hat{\mathcal{B}}$. Meanwhile, $\rho_{\bar{\ell},\ell}$ fluctuates near 0.4 throughout training (see Fig. 6a). In Lunar-Lander, Fig. 6b shows that MPTS can keep the leading trend in both average and CVaR performance. However, DRM and GDRM not only fall behind others in the average returns but also receive the lowest $\text{CVaR}_{0.9}$ returns, completely failing in robust optimization. The simple worst-case task selection or reweight might deteriorate overall performance when some challenging tasks are not available to be successfully resolved while being sampled with a high frequency. Instead, MPTS enlarges the search scope and selects a subset based on the rule hybrid of the worst-case and the uncertainty, which helps avoid over-optimizing the finite worst MDPs. Regarding $\rho_{\bar{\ell},\ell}$, it quickly peaks above 0.6 and stabilizes around 0.3 over iterations. This is in accordance with prior evidence; the selected tasks around the convergence iteration might concentrate and weaken the discrimination of task difficulties.
@@
In Lunar-Lander, $\bm{\tau}\in\mathbb{R}^{+}$ means the main engine strength of the rocket. We shift the DR-training identifier’s interval from $\bm{\tau}\in[4.0,20.0]$ to $\bm{\tau}\in[1.0,4.0)\cup(20.0,23.0]$ and test DR-trained policies on these OOD MDPs. Compared with the tested ID MDP returns (see the middle of Fig. 6g), all methods encounter sharp performance decline in the hard OOD tasks and retain faint return fluctuations in the easy OOD tasks $\bm{\tau}\in(20.0,23.0]$. DRM-trained policies exhibit more deviations and are prone to instability in the deployment of easy tasks. However, MPTS still dominates returns over others in each of the hard OOD tasks, exhibiting strong adaptation robustness in difficult cases.
@@ 3.4 MPTS also reserves the potential of robust PEFT
Figure 7: Testing Classification Results after Prompt-Tuning on ImageNet. a-d. Shown are testing $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average accuracies with the prompt-tuned machine learner on ID and OOD datasets. e. During prompt-tuning ImageNet, we report the memory cost and clock time relative to ERM for all methods.
@@
Fig. 7a-d indicates that MPTS can consistently outperform other baselines in average and different CVaR accuracies on ID and OOD testing datasets. MPTS beats ERM by 0.82%-3.11% accuracy increase in terms of $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$ and $\text{CVaR}_{0.5}$ indicators (see Table 6). MPTS is more advantageous on OOD datasets than ImageNet results Overall, on 4/5 datasets except ImageNet-A, DRM is nearly SOTA secondary to MPTS on CVaR accuracies but reaches ERM comparable performance in average accuracies. GDRM’s performance varies with $\alpha$ levels and shows marginal gains compared with ERM. Like the few-shot image classification case, DRM sacrifices the computational and memory efficiency for specific robustness improvement as reported in Fig. 7e. Even though MPTS shares a similar optimization purpose with DRM, the risk predictive module and simulation over larger batches empower MPTS with more task exploration advantage at negligible computational cost, resulting in a more optimal robust machine learner.
@@ 4 Discussion
Rapid adaptation to novel scenarios with any-shot examples has been identified as an indispensable capability in artificial general intelligence. While the neural scaling laws stress the importance of the scale of task datasets, practical considerations, such as safety issues and limited data and computational budgets, necessitate a closer focus on adaptation robustness and learning efficiency. This study takes a closer look at the optimization process of any-shot adaptation through generative modeling and presents MPTS as a versatile module for robust active task sampling.
@@
First of all, this work reports a vital finding that it is theoretically feasible to predict the optimization outcome in a rough granularity for the online scoring difficulty of any task for the machine learner to adapt. Extensive experimental results demonstrate that MPTS can (i) amortize any-shot adaptation evaluation for robustness purposes, (ii) significantly enhance adaptation robustness across diverse tasks, including pattern recognition and decision-making, and (iii) accelerate the learning process while promoting exploration in the task space under specific scenarios.
@@
This work’s main contribution is a novel framework of task sampling to promote robustness without sacrificing learning efficiency. The large-scale experimental results show the promise of risk predictive models in surrogating adaptation performance evaluation, which can be integrated with more task selection rules in developing the foundation models. However, in the presence of multi-modal signals, seeking risk learners to process tasks deserves more effort in the future.
@@ 5 Methods
In alignment with the realistic necessities, this work focuses on robust adaptation while securing learning efficiency, such as circumventing expensive evaluation to allow for real-world adaptation. Such a purpose facilitates the birth of MPTS. As previously mentioned, the framework is agnostic to specific task episodic learning methods; hence, we leave out any-shot learning details. As noticed in Fig. 1a, several roles are involved in the optimization: (1) the adaptive machine learner, e.g., foundation models or generalist policies, learns to learn in an any-shot way given some optimizers; (2) the risk learner as a critic evaluates and forecasts the task-specific adaptation risk; (3) the task sampler as an actor works for screening task batches for next iteration. These components operate in a cycle until convergence.
@@ 5.1 Generative Modeling Risk Functions and Posterior Inference
Here, we design the sampling strategy through the lens of risk landscapes and pay more attention to learning optimization outcome datasets $\{H_{t}\}_{t=1}^{T}$. To characterize the adaptation risk during batch optimization, we introduce the latent variable $\bm{z}$ to summarize episodic information and present a versatile deep generative model in Eq. (9).
$\begin{split}p(H_{0:T},\bm{z}_{0:T}|\bm{\theta}_{0:T})=p(\bm{z}_{0})\prod_{t=0}^{T}p_{\bm{\psi}}(H_{t}|\bm{z}_{t};\bm{\theta}_{t})\prod_{t=0}^{T-1}p(\bm{z}_{t+1}|\bm{z}_{t})\end{split}$ (9)
Within a Bayesian framework, we approximate the underlying function distribution with the latent variable, and the posterior $p(\bm{z}_{t}|H_{t})$ over $\bm{z}$ summarizes the historical risk information and accounts for uncertainty in distributions. The following writes the form of $p(\bm{z}_{t}|H_{t})$ according to Bayes rule 53:
@@ 5.2 Task Sampling Strategy Design
In robust active task sampling, existing strategies evaluate task batches to score their difficulties in adaptation and then prioritize challenging subsets for optimization 20, 16. Besides the expensive evaluation cost, these strategies are weak in the efficient exploration of the task space.
As Theorem 1 has established the theoretical foundation of amortized evaluation, this necessitates the development of the risk learner from cumulated risk histories. With the model predictive results as amortized evaluation, specific rules can be incorporated into the acquisition function for effective sampling.
@@ 5.3 Sequentially Optimize the Adaptive Machine Learner
Here, we take the finetuning pretrained models to downstream tasks 61 as an instantiation. In this case, each data point $[\bm{x},\bm{y}]$ can be viewed as a task with either its embedding $\bm{\tau}$ or $\bm{x}$ as the task identifier. Then the model update rule can be:
@@ Appendix A Quick Guideline to MPTS
The involvement of the risk learner inevitably brings extra computational overhead in optimization. However, the risk learner used in this work is lightweight with the model complexity $\mathcal{O}(|\bm{\phi}|+|\bm{\psi}|)<<\mathcal{O}(|\bm{\theta}|)$. We can roughly estimate these extra computations that arise from the predictive model as $\mathcal{O}\big{(}(|\bm{\phi}|+|\bm{\psi}|)T_{\text{MPTS}}\big{)}$ throughout the training phase. Moreover, the computational and task evaluation complexities of different methods are estimated in Table 2. Compared with DRM, MPTS retains more computational and task efficiency when the filtering ratio $\alpha_{\text{train}}$ is high, and the machine learner $\bm{\theta}$ is largely given similar convergence iteration steps. However, exactly evaluating task difficulties from either losses or gradients brings extra computational and annotation overhead.
@@
computation $\mathcal{O}(|\bm{\theta}|T_{\text{ERM}})$ $\mathcal{O}(\frac{1}{1-\alpha_{\text{train}}}|\bm{\theta}|T_{\text{DRM}})$ $\mathcal{O}(|\bm{\theta}|T_{\text{GDRM}})$ $\mathcal{O}\big{(}(|\bm{\phi}|+|\bm{\psi}|+|\bm{\theta}|)T_{\text{MPTS}}\big{)}$
task eval $\mathcal{O}(\mathcal{B}T_{\text{ERM}})$ $\mathcal{O}(\frac{\mathcal{B}}{1-\alpha_{\text{train}}}T_{\text{DRM}})$ $\mathcal{O}(\mathcal{B}T_{\text{GDRM}})$ $\mathcal{O}(\mathcal{B}T_{\text{MPTS}})$
@@ B.2 Dataset Curation and Task-Level Robustness
Another strategy to evaluate the machine learner’s robustness is the performance in OOD tasks. This refers to the case when the training and the testing task distributions are different. Particularly, in DR and prompt-tuning scenarios, we also use the OOD tasks that never appear in the training task distribution to test the trained policy, and this setup corresponds to domain generalization, a type of substantial distribution shift 13.
@@ C.2 Tasks with Implicit Identifiers
Nevertheless, in several scenarios, it is intractable to access the explicit task identifier. For example, in few-shot image classification, the task information is just the coupled support and query dataset $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}$. Similarly, in PEFT for LLMs, the task can be in the form of the QA pair $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{Q}$. There is no explicit representation method, such as $\bm{\tau}$, for these tasks, which brings difficulty in building up the risk learner. Retaining the prior notation, the episodic task batch can be written as $\hat{H}_{t}=\left\{\bm{\theta}_{t},\left({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{\tau}_{t,i}},\mathcal{D}_{\bm{\tau}_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$, where $\bm{\tau}$ of our interest is unobservable. Some experiments in this work, such as few-shot image classification and PEFT, encounter such embarrassment.
@@
Fortunately, there exist pretrained models that enable the task representation to be generalizable to downstream tasks. For example, in the N-way K-shot image classification, the task is in the form of support image-label pairs and the query images and the goal is to assign labels to the query images from the support dataset. With the help of CLIP models 1, for a fixed task in the form of $\mathcal{D}_{\tau}$, we can access a N vectors $\{\bm{z}_{i}\}_{i=1}^{N}$ by inputing the set of text-form classes $\{\bm{c}_{i}\}_{i=1}^{N}$ extracted from the support dataset $\mathcal{D}_{\tau}^{S}$, i.e., $\text{CLIP}(\{\bm{c}_{i}\}_{i=1}^{N})=\{\bm{c}_{i}\}_{i=1}^{N}:=\bm{\tau}$. As a result, we can obtain $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$ conditioned on $\bm{\theta}_{t}$ for feasible task risk functional prosterior inference. This helps our approach to circumvent the unavailability of exact task identifiers. And it is plausible for the risk predictive model to optimize in learning $p(\ell|\bm{\tau},H_{1:t})$. It is worth noting that this case still prefers lightweight models for identifier inference, and the text encoder of CLIP well satisfies this requirement and can be used in the N-way K-shot image classification. Details on specific task identifier inference modules can be found in Section E and F.

### v3

@@ 3.3 MPTS retains multi-faced advantages beyond robustness in zero-shot continuous control
Figure 6: DR Results on Ergo-Reacher and Lunar-Lander (7 Runs). a. In Ergo-Reacher, the $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average cumulative returns on validation MDPs are reported together with the risk learner’s PCC curve during DR training. b. In Lunar-Lander, the cumulative returns on validation MDPs are illustrated together with the risk learner’s PCC curve during DR training. c. We test the DR-trained policies on the fixed MDP set and report the $\text{CVaR}_{\alpha}$ cumulative returns. d. The returns averaged over DR-testing MDPs are illustrated. e. The required runtime is computed for all methods on Lunar-Lander. f. In Lunar-Lander, shown are frequencies of sampled identifiers using MPTS during DR training. g. In Lunar-Lander, we test the trained policies in both in-distribution (ID) domains and out-of-distribution (OOD) domains to report each task’s average returns.
MPTS dominates the overall performance in DR training.
The Ergo-Reacher curves oppose general Meta-RL findings; the use of robust optimization, such as MPTS and DRM, improves both the average and CVaR performance in Fig. 6a. One probable reason is that MPTS benefits from extra task exploration with the relatively larger $\hat{\mathcal{B}}$. Meanwhile, $\rho_{\bar{\ell},\ell}$ fluctuates near 0.4 throughout training (see Fig. 6a). In Lunar-Lander, Fig. 6b shows that MPTS can keep the leading trend in both average and CVaR performance. However, DRM and GDRM not only fall behind others in the average returns but also receive the lowest $\text{CVaR}_{0.9}$ returns, completely failing in robust optimization. The simple worst-case task selection or reweighted mechanism might deteriorate overall performance when some challenging tasks unavailable to be successfully resolved are yet sampled with a high frequency. Instead, MPTS enlarges the search scope and selects a subset based on the rule hybrid of the worst-case and the uncertainty, which avoids over-optimizing the finite worst MDPs. Regarding $\rho_{\bar{\ell},\ell}$, it quickly peaks above 0.6 and stabilizes around 0.3 over iterations. This is in accordance with prior evidence; the selected tasks around the convergence iteration might concentrate and weaken the discrimination of task difficulties.
@@
In Lunar-Lander, $\bm{\tau}\in\mathbb{R}^{+}$ means the main engine strength of the rocket. We shift the DR-training identifier’s interval from $\bm{\tau}\in[4.0,20.0]$ to $\bm{\tau}\in[1.0,4.0)\cup(20.0,23.0]$ and test DR-trained policies on these OOD MDPs. Compared with the tested ID MDP returns (see the middle of Fig. 6g), all methods encounter sharp performance decline in the hard OOD tasks and retain faint return fluctuations in the easy OOD tasks $\bm{\tau}\in(20.0,23.0]$. DRM-trained policies exhibit more deviations and fail to advance generalization to easier tasks. However, MPTS still dominates returns over others in the hard OOD tasks, exhibiting strong adaptation in difficult cases.
@@ 3.4 MPTS also reserves the potential of robust PEFT
Figure 7: Testing Classification Results after Prompt-Tuning on ImageNet (3 Runs in Average). a-d. Shown are testing $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average accuracies with the prompt-tuned machine learner on ID and OOD datasets. e. During prompt-tuning ImageNet, we report the memory cost and clock time relative to ERM for all methods.
@@
In Fig. 7a-d, MPTS consistently outperforms other baselines in average and different CVaR accuracies on ID and OOD testing datasets. MPTS beats ERM by 0.82%-3.11% accuracy increase in terms of $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$ and $\text{CVaR}_{0.5}$ indicators (see Table 6). MPTS is more advantageous on OOD datasets than ImageNet results. On 4/5 datasets except ImageNet-A, DRM is nearly SOTA secondary to MPTS on CVaR accuracies but reaches ERM comparable performance in average accuracies. GDRM’s performance varies with $\alpha$ levels and shows marginal gains compared with ERM. Like the few-shot image classification case, DRM again sacrifices the computational and memory efficiency for specific robustness improvement as reported in Fig. 7e. Even though MPTS shares a similar optimization purpose with DRM, the risk predictive module and simulation over larger batches empower MPTS with more task exploration power at negligible computational cost, resulting in a more optimal robust machine learner.
@@ 4 Discussion
Rapid adaptation to novel scenarios with any-shot examples has been identified as an indispensable capability in artificial general intelligence. While the neural scaling laws stress the importance of the scale of task datasets, practical considerations, such as safety issues and limited annotation and computational budgets, necessitate a closer focus on adaptation robustness and learning efficiency. This study takes a closer look at the optimization process of any-shot adaptation through generative modeling and presents MPTS as a versatile module for robust active task sampling.
@@
First of all, this work reports a vital finding that it is theoretically feasible to predict the optimization outcome in a rough granularity for the online scoring difficulty of any task for the machine learner to adapt. Extensive experimental results demonstrate that MPTS can (i) amortize any-shot adaptation evaluation for robustness purposes and promote exploration in the task space, (ii) significantly enhance adaptation robustness across diverse tasks, including pattern recognition and decision-making, and (iii) accelerate the learning process under scenarios like few-shot regression and DR.
@@
This work’s main contribution is a novel framework of task sampling to promote robustness without sacrificing learning efficiency. The large-scale experimental results show the promise of risk predictive models in surrogating adaptation performance evaluation, which can be integrated with more task selection rules in developing the foundation models. However, more effort is needed to scale MPTS to training or tuning foundation models with more complicated multimodal signals.
@@ 5 Methods
In alignment with the realistic necessities, this work focuses on robust adaptation while securing learning efficiency, such as circumventing expensive evaluation to allow for real-world adaptation. Such a purpose facilitates the birth of MPTS. As previously mentioned, the framework is agnostic to task episodic learning methods; hence, we leave out any-shot learning details. As noticed in Fig. 1a, several roles are involved in the optimization: (1) the adaptive machine learner, e.g., foundation models or generalist policies, learns to learn in an any-shot way given some optimizers; (2) the risk learner as a critic evaluates and forecasts the task-specific adaptation risk; (3) the task sampler as an actor works for screening task batches for next iteration. These components participate in episodic learning until convergence.
@@ 5.1 Generative Modeling Risk Functions and Posterior Inference
Here, we design the sampling strategy through the lens of risk landscapes and pay more attention to datasets of learning optimization outcome $\{H_{t}\}_{t=1}^{T}$. To characterize the adaptation risk during batch optimization, we introduce the latent variable $\bm{z}_{t}$ to summarize episodic information and present a versatile deep generative model as:
$\begin{split}p(H_{0:T},\bm{z}_{0:T}|\bm{\theta}_{0:T})=p(\bm{z}_{0})\prod_{t=0}^{T}p_{\bm{\psi}}(H_{t}|\bm{z}_{t};\bm{\theta}_{t})\prod_{t=0}^{T-1}p(\bm{z}_{t+1}|\bm{z}_{t}).\end{split}$ (9)
Within a Bayesian framework, we approximate the underlying function distribution with the latent variable, and the posterior $p(\bm{z}_{t}|H_{t})$ summarizes the historical risk information and accounts for uncertainty in distributions. The following writes the form of $p(\bm{z}_{t}|H_{t})$ according to the Bayes rule 53:
@@ 5.2 Task Sampling Strategy Design
In robust active task sampling, existing strategies evaluate task batches to rank their difficulties in adaptation and then prioritize challenging subsets for optimization 20, 16. Besides the expensive evaluation cost, these strategies are weak in the efficient exploration of the task space.
As Theorem 1 has established the theoretical foundation of amortized evaluation, this necessitates the development of the risk learner from cumulated risk histories. With the model predictive results as amortized evaluation, specific rules can be incorporated into the acquisition function for active sampling.
@@ 5.3 Sequentially Optimize the Adaptive Machine Learner
Here, we take finetuning pretrained models to downstream tasks 61 as an instantiation. In this case, each data point $[\bm{x},\bm{y}]$ can be viewed as a task with either its embedding $\bm{\tau}$ or $\bm{x}$ as the task identifier. Then the model update rule can be:
@@ Appendix A Quick Guideline to MPTS
The involvement of the risk learner inevitably brings extra computational overhead in optimization. However, the risk learner used in this work is lightweight with the model complexity $\mathcal{O}(|\bm{\phi}|+|\bm{\psi}|)<<\mathcal{O}(|\bm{\theta}|)$. We can roughly estimate these extra computations that arise from the predictive model as $\mathcal{O}\big{(}(|\bm{\phi}|+|\bm{\psi}|)T_{\text{MPTS}}\big{)}$ throughout the training phase. Moreover, the computational and task evaluation complexities of different methods are estimated in Table 2. Compared with DRM, MPTS retains more computational and task efficiency when the filtering ratio $\hat{\alpha}$ is high, and the machine learner $\bm{\theta}$ is largely given similar convergence iteration steps. However, exactly evaluating task difficulties from either losses or gradients brings extra computational and annotation overhead.
@@
computation $\mathcal{O}(|\bm{\theta}|T_{\text{ERM}})$ $\mathcal{O}(\frac{1}{1-\hat{\alpha}}|\bm{\theta}|T_{\text{DRM}})$ $\mathcal{O}(|\bm{\theta}|T_{\text{GDRM}})$ $\mathcal{O}\big{(}(|\bm{\phi}|+|\bm{\psi}|+|\bm{\theta}|)T_{\text{MPTS}}\big{)}$
task eval $\mathcal{O}(\mathcal{B}T_{\text{ERM}})$ $\mathcal{O}(\frac{\mathcal{B}}{1-\hat{\alpha}}T_{\text{DRM}})$ $\mathcal{O}(\mathcal{B}T_{\text{GDRM}})$ $\mathcal{O}(\mathcal{B}T_{\text{MPTS}})$
@@ B.2 Dataset Curation and Task-Level Robustness
Another strategy to evaluate the machine learner’s robustness is the performance in OOD tasks. This refers to the case when the training and the testing task distributions are different. Particularly, in DR and prompt-tuning scenarios, we also use the OOD tasks that never appear in the training task distribution to test the trained policy, and this setup corresponds to domain generalization, a type of substantial distribution shift 11.
@@ C.2 Tasks with Implicit Identifiers
Nevertheless, in several scenarios, it is intractable to access the explicit task identifier. For example, in few-shot image classification, the task information is just the coupled support and query dataset $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}$. Similarly, in PEFT for LLMs, the task can be in the form of the QA pair $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{Q}$. There is no explicit representation method, such as $\bm{\tau}$, for these tasks, which brings difficulty in building up the risk learner. Retaining the prior notation, the episodic task batch can be written as $\hat{H}_{t}=\left\{\bm{\theta}_{t},\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$, where $\bm{\tau}$ of our interest is unobservable. Some experiments in this work, such as few-shot image classification and PEFT, encounter such embarrassment.
@@
Fortunately, there exist pretrained models that enable the task representation to be generalizable to downstream tasks. For example, in the N-way K-shot image classification, the task is in the form of support image-label pairs and the query images and the goal is to assign labels to the query images from the support dataset. With the help of CLIP models 1, for a fixed task in the form of $\mathcal{D}_{\tau}$, we can access a N vectors $\{\bm{z}_{i}\}_{i=1}^{N}$ by inputing the set of text-form classes $\{\mathcal{C}_{i}\}_{i=1}^{N}$ extracted from the support dataset $\mathcal{D}_{\tau}^{S}$, i.e., $\text{CLIP}(\{\mathcal{C}_{i}\}_{i=1}^{N})=[\text{CLIP}_{\text{text}}(\mathcal{C}_{1}),\dots,\text{CLIP}_{\text{text}}(\mathcal{C}_{K})]:=\bm{\tau}$. As a result, we can obtain $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$ conditioned on $\bm{\theta}_{t}$ for feasible task risk functional prosterior inference. This helps our approach to circumvent the unavailability of exact task identifiers. And it is plausible for the risk predictive model to optimize in learning $p(\ell|\bm{\tau},H_{1:t})$. It is worth noting that this case still prefers lightweight models for identifier inference, and the text encoder of CLIP well satisfies this requirement and can be used in the N-way K-shot image classification. Details on specific task identifier inference modules can be found in Section E and F.